OpenStudy (coconut):

What is the 7th term of the geometric sequence where a1 = 1024 and a3 = 64?

OpenStudy (anonymous):

do you know the formula ? :)

OpenStudy (coconut):

no sorry :(

Parth (parthkohli):

You are dividing by 4 every time @Kreshnik

Parth (parthkohli):

a4 = 16, a5 = 4, a6 = 1, a7 = 1/4 = 0.25

Parth (parthkohli):

0.25 is the 7th term.

Parth (parthkohli):

64/4 = 16, 16/4 = 4, 4/4 = 1, 1/4 = 0.25

Parth (parthkohli):

Do you get it?

OpenStudy (anonymous):

@coconut \[\LARGE a_n=a_1\cdot r^{n-1}\] let's substitute: \[\LARGE a_3=1024\cdot r^{3-1}\] \[\LARGE 64=1024\cdot r^2\] \[\LARGE r^2=\frac{64}{1024}\] \[\LARGE r^2=\frac{16}{256 }\] \[\LARGE r^2=\frac{8}{128 }\] \[\LARGE r^2=\frac{8}{128 }\] \[\LARGE r^2=\frac{1}{16 }\] \[\LARGE r^2=\frac{1}{4^2 }\] \[\LARGE r^2=\left(\frac{1}{4 }\right)^2\] \[\LARGE r=\frac14 \]

OpenStudy (anonymous):

now you have: \[\LARGE a_7=a_1\cdot r^{7-1}\] \[\LARGE a_7=1024\cdot \left(\frac14 \right)^{6}\] go for it...

OpenStudy (anonymous):

@ParthKohli Yes Indeed 7th term is 0.25 ^^ Well done.